OpenStudy (anonymous):

Find the value of x: -28 = -196x x = 1/7 x = -7 x = 7 x = -1/7

OpenStudy (anonymous):

x=1/7

OpenStudy (idealist10):

-28=-196x 28=196x x=28/196 x=1/7 Answer: 1/7